OpenStudy (ambermarie151):

solve this quadratic equation by completing the square x^2+6x =18

jhonyy9 (jhonyy9):

x^2 +6x =18 add to both sides 9 x^2 +6x +9 = 27 (x+3)^2 = 27 x+3 = +/- 3sqrt3 x_1,2 = -3 +/- 3sqrt3 hope helped